OpenStudy (anonymous):

what is tan 60 using the unit circle ?

OpenStudy (anonymous):

tan = sin/cos sin 60 = sqrt(3)/2 cos 60 = 1/2 (sqrt(3)/2)/(1/2) = sqrt(3)

OpenStudy (anonymous):

thanks bro